Find a if the point (3,a) is on the line that passes through (-2,7) and (5,-3).
Answer by Cromlix(4381)

Hi there,
First make up the equation of
the line between (-2,7) and(5,-3)
Gradient = y2 - y1/x2 - x1
Gradient = -3-7/5-(-2)
Gradient = -10/7
Using equation of line formula:
y - b = m(x - a)
using m = -10/7 and (a,b)= (-2,7)
y - 7 = -10/7(x - (-2))
y - 7 = -10/7(x + 2)
y - 7 = -10/7x - 20/7
y = -10/7x - 20/7 + 49/7 (7)
y = -10/7x + 29/7
This is your equation.......
............................
Now substitute the points coordinates
into the equation:
y = -10/7x + 29/7
y = -10/7(3) + 29/7
y = -30/7 + 29/7
So, y = -1/7
Point coords (3, -1/7)
Hope this helps:-)